# PUP 3D-GS: Principled Uncertainty Pruning for 3D Gaussian Splatting

### v1

@@ 4.3 Patch-wise Uncertainty
As mentioned in Section 4.1, the computation of the Hessian over the entire scene requires a sum over all per-pixel Fisher approximations of the reconstructed input views. However, we empirically observe that our sensitivity scores computed over image patches maintain a high correlation with our sensitivity scores computed over individual pixels. Specifically, we compute the Fisher approximation on image patches by applying an average pool over each $n\times n$ block of pixels in the rendered images, then compute the Fisher approximation on each pixel of the downsampled image. To obtain the scene-level Hessian we sum all patch-wise Fisher approximations over all images. In our experiments, we compute Fisher approximations over $8\times 8$ image patches, but Figure 3 illustrates that the correlation with per-pixel scores is maintained even with larger patches of size $64\times 64$. We further ablate our choice of patch size in Appendix A.3.
@@ 5.1 Datasets
We evaluate our approach on the same challenging, real world scenes as 3D-GS [9]. All nine scenes from the MiP-NeRF 360 dataset [1], which consists of five outdoor and four indoor scenes that each contain a complex central object or viewing area and a detailed background, are used. Two outdoor scenes, truck and train, are taken from the Tanks & Temples dataset [10], and two indoor scenes, drjohnson and playroom, are taken from the Deep Blending dataset [6]. For consistency, we use the COLMAP [21] camera pose estimates that are provided by the creators of the datasets.
@@ 5.2 Implementation Details
Our method, PUP 3D-GS, can be applied to any 3D-GS model. We implement the Hessian computation on the original 3D-GS codebase [9], then adapt the pruning and refining framework from LightGaussian [3] to accommodate our uncertainty estimate and multi-round pruning method. For a fair comparison, we run LightGaussian and our pipeline on the same pretrained 3D-GS scenes. Rendering speeds are collected using a Nvidia RTXA5000 GPU in frames per second (FPS).
@@ 5.3.1 Analysis of PUP 3D-GS Pipeline
The efficacy of our PUP 3D-GS pipeline is demonstrated by Table 1, where we record the mean of each metric over the Mip-NeRF 360 dataset. Pruning $66\%$ of Gaussians from the original 3D-GS model nearly doubles its FPS with only a small degradation of PSNR, SSIM, and LPIPS. Fine-tuning the pruned model for 5,000 iterations recovers the image quality metrics while only slightly decreasing rendering speed. Pruning $66\%$ of the remaining Gaussians, such that a total of $88.44\%$ of Gaussians are pruned from the base 3D-GS scene, causes a larger drop in the image quality metrics. However, we recover most of this degradation by fine-tuning this significantly smaller model for another 5,000 iterations. Rendering speed is nearly tripled.
@@ 5.3.2 Comparison with LightGaussian
Table 2 compares our $88.44\%$ pruning results against LightGaussian’s using the per-dataset mean of each metric. For our method, we run the full two step prune-refine pipeline described in Section 5.3.1, where we perform two rounds of $66\%$ pruning using our spatial uncertainty estimate with 5,000 fine-tuning iterations each. In LightGaussian, we run their one-round pipeline to prune $88.44\%$ of Gaussians using their global significance score and then fine-tune for 10,000 iterations. In LightGaussian (2×2$\times$2 ×), we use their global significance score in our two step prune-refine pipeline.
We show that our spatial uncertainty estimate outperforms both configurations of LightGaussian across nearly every metric. The lone exception is PSNR on the Tanks & Temples dataset, where we are slightly outperformed by LightGaussian by less than 0.06 points.
Table 2: Comparison of our PUP 3D-GS pipeline against LightGaussian [3]. For a fairness, we record the results of the original LightGaussian pipeline which uses a single prune-refine step and the results of applying our multi-round pruning to the LightGaussian pipeline, LightGaussian (2$\times$), where we run their prune-refine step twice. The original LightGaussian pipeline prunes to $88.44\%$ once, while our pipeline and LightGaussian (2$\times$) prune $66\%$ twice for a total of $88.44\%$: the final sizes are identical. PUP 3D-GS increases the rendering speed of 3D-GS by 2.65×2.65$\times$2.65 × on average.
@@
MipNerf-360 3D-GS 27.47 0.812 0.221 95.59 746.46
LightGaussian 26.25 0.763 0.298 177.4 86.29
LightGaussian (2$\times$) 26.18 0.761 0.304 191.36 86.29
Ours 26.83 0.792 0.268 244.07 86.29
Tanks & Temples 3D-GS 23.77 0.845 0.1777 147.97 433.24
LightGaussian 23.09 0.797 0.2603 370.42 50.08
LightGaussian (2$\times$) 23.03 0.793 0.2663 320.12 50.08
Ours 23.03 0.807 0.2453 417.98 50.08
Deep Blending 3D-GS 28.98 0.8816 0.2856 105.61 699.19
LightGaussian 28.35 0.8676 0.3333 240.17 80.83
LightGaussian (2$\times$) 28.11 0.8647 0.3373 210.12 80.83
Ours 28.61 0.8812 0.3053 296.08 80.83
Figure 4: Visual comparison after two rounds of prune-refine using our method and LightGaussian’s method. Top: bonsai from the Mip-Nerf 360 dataset. Middle: room from the Mip-Nerf 360 dataset. Bottom: train from the Tanks & Temples dataset. Additional visual results are in Appendix A.2
In Figure 4, we report qualitative results on one scene taken from each dataset. The magnified regions demonstrate that our method retains many fine foreground details that are not preserved by LightGaussian. Error residual images with respect to the ground truth and original 3D-GS scenes are available in Appendix A.6 and Appendix A.7; visualizations of other scenes are provided in Appendix A.2.
We speculate that the visibility score in LightGaussian’s importance sampling heuristic is biased towards retaining larger Gaussian because they have a higher probability of being visible in more pixels across the training views. In Figure 5, we plot the distribution of the log determinant of the Gaussian covariances over the scenes visualized in Figure 4. The distribution clearly skews more heavily towards larger Gaussians in the scenes pruned with LightGaussian than our pipeline.
Figure 5: Histograms showing the distribution of Gaussians over the log of their volumes for the bonsai, room, and train scenes after two rounds of prune-refine with our PUP 3D-GS pipeline and LightGaussian’s pipeline (LG). LightGaussian consistently retains larger Gaussians than ours.
@@ 5.3.3 Vectree Quantization
We apply the Vector Quantization implementation found in the LightGaussian framework to further compress our pruned models. Our final average model sizes are $16.76$MB for Mip-NeRF 360, $9.87$MB for Tanks & Temples, and $15.64$MB on Deep Blending. In other words, the unpruned 3D-GS scenes are reduced in size by 44.5×44.5$\times$44.5 ×, 43.9×43.9$\times$43.9 ×, and 44.7×44.7$\times$44.7 ×, respectively.
Table 3: Vectree Quantization comparison with LightGaussian after pruning to $88.44\%$. Similar to Table 2, we report quantization on the original one-round LightGaussian method as well LightGaussian with two rounds of prune-refine. PUP 3D-GS outperforms LightGaussian across all metrics.
@@
Mip-NeRF 360 3D-GS 27.47 0.8123 0.2216 95.59 746.46
LightGaussian 24.61 0.7272 0.3313 217.64 16.8
LightGaussian (2$\times$) 24.58 0.7252 0.3318 199.1 16.81
Ours 25.05 0.7599 0.2963 257.57 16.76
Tanks & Temples 3D-GS 23.77 0.8458 0.1777 147.97 433.24
LightGaussian 21.80 0.7666 0.2897 412.94 9.91
LightGaussian (2$\times$) 21.80 0.7668 0.2896 367.87 9.92
Ours 21.81 0.7837 0.2673 428.59 9.87
Deep Blending 3D-GS 28.98 0.8816 0.2859 105.61 699.19
LightGaussian 27.66 0.8564 0.3467 238.94 15.65
LightGaussian (2$\times$) 27.37 0.8526 0.3474 254.64 15.65
Ours 27.79 0.8700 0.3146 351.72 15.64
@@ 6 Ablations
6.1 Multi-Round vs. One-Round Pipeline
A core component of our pipeline is running prune-refine over multiple rounds. In Table 4, we compare the mean image quality and FPS over the Mip-NeRF 360 scenes when we prune using our two step prune-refine method against when we prune 88.44% of Gaussians in one round. Multi-round pruning produces better measurements across all metrics.
Table 4: One step of pruning $88.44\%$ of Gaussians from 3D-GS and then fine-tuning for 10,000 iterations, versus two steps of pruning $66\%$ of Gaussians and finetuning for 5,000 iterations each. In both cases, $88.44\%$ Gaussians are pruned and the model is fine-tuned for 10,000 total steps. Our multi-round approach outperforms the one-round approach across all metrics.
@@
Prune $88.44\%+10$K Finetune 26.51 0.7864 0.2731 221.34 86.29
(Prune $66\%+5$K Finetune) ×$\times 2$ 26.83 0.7920 0.2688 244.07 86.29
@@
Another component of our method is restricting our sensitivity pruning score to only the mean and scale parameters for each Gaussian. Table 5 shows that image quality decreases if we use the RGB color parameters of the Gaussians to compute our sensitivity score instead. We speculate that the lower performance of the RGB sensitivity score is due to its similarity with LightGaussian’s visibility score – the change in the L2 error over the training images with respect to perturbations in the RGB value of a Gaussian correlates with its visibility across the training views.
Table 5: Results from our two step prune-refine pipeline when computing sensitivity with spatial and RGB color parameters. The spatial score produces higher fidelity renderings than the color score.
Methods Mip-NeRF 360 Bicycle Scene
@@
3D-GS 25.09 0.7467 0.2442 50.23 1345.58
RGB Sensitivity 24.73 0.7155 0.2963 160.81 155.55
Spatial Sensitivity 25.00 0.7270 0.2870 139.67 155.55
@@ 7 Limitations
A limitation of our approach is that it assumes that the scene is converged because the Fisher approximation requires a small L1 residual to be accurate. We find that a small residual is preserved even after removing $66\%$ of the Gaussians with prune-refine, allowing us to perform a second round, as shown in Section 4.4. Our Fisher approximation also involves computational overhead that is reduced by performing it in the patch-wise manner described in Section 4.3. We are optimistic that more sophisticated techniques like optimized CUDA kernels and data parallelization can further improve the efficiency of our approximation.
@@ 8 Conclusion
In this work, we propose PUP 3D-GS: a new post-hoc pipeline for pruning pretrained 3D-GS models without changing their training pipelines. It uses a novel, mathematically principled approach for choosing which Gaussians to prune by assessing their sensitivity to the reconstruction error over the training views. This sensitivity pruning score is derived from a computationally tractable Fisher approximation of the Hessian over that reconstruction error. We use our sensitivity score to prune the reconstructed scene, then fine-tune the remaining Gaussians in the model over multiple rounds. We observe that PUP 3D-GS improves the rendering speed of 3D-GS by 2.65×2.65$\times$2.65 × on average, while retaining more salient foreground information than existing techniques and achieving higher image quality metrics on the MiP-NeRF 360, Tanks & Temples, and Deep Blending dataset scenes.
@@ Appendix A Appendix
A.1 Scene Evaluations
PSNR, SSIM, LPIPS, and FPS for each scene from the Mip-NeRF 360, Tanks&Temples, and Deep Blending datasets that was used in 3D-GS [9]. Note that the sizes of the pruned scenes in this section are identical because exactly $88.44\%$ of Gaussians were removed from each of them using our two step prune-refine method.

### v3

@@ 4.3 Patch-wise Uncertainty
The computation of the Hessian over the entire scene requires a sum over all per-pixel Fisher approximations of the reconstructed input views. However, we empirically observe that sensitivity scores computed over image patches are highly correlated with those computed over individual pixels. Specifically, we compute the Fisher information approximation on image patches by rendering images at lower resolution and then computing the Fisher approximation on each of their pixels. Then, we obtain the scene-level Hessian by summing the patch-wise Fisher approximations over all views. We use $4\times 4$ image patches in our experiments and ablate our choice of patch size in Appendix A.2.
@@ 4.4 Multi-Round Pipeline
Figure 3: The average metrics over all scenes in the Mip-NeRF 360 dataset after pruning with different percentages in our two-round pipeline. PSNR and SSIM decrease and LPIPS and FPS increase with more pruning. Per-round percentages are selected in $5\%$ intervals and the model is fine-tuned for $5,000$ iterations in each round. The red dots at $(80\%,50\%)$ represent our percentages for $90\%$ compression.
@@ 5.1 Datasets
We evaluate our approach on the same challenging, real world scenes as 3D-GS [11]. All nine scenes from the Mip-NeRF 360 dataset [1], which consists of five outdoor and four indoor scenes that each contain a complex central object or viewing area and a detailed background, are used. Two outdoor scenes, truck and train, are taken from the Tanks & Temples dataset [13], and two indoor scenes, drjohnson and playroom, are taken from the Deep Blending dataset [8]. For consistency, we use the COLMAP [28] camera pose estimates that were provided by the creators of the datasets.
@@ 5.2 Implementation Details
Our method, PUP 3D-GS, can be applied to any 3D-GS model. We implement the Hessian computation as a CUDA kernel in the original 3D-GS codebase [11], then adapt the pruning and refining framework from LightGaussian [4] to accommodate our uncertainty estimate and multi-round pruning method. Computing pruning scores across the entire training view set takes seconds. For a fair comparison, we run LightGaussian and our pipeline on the same pretrained 3D-GS scenes. Rendering speeds are collected using a Nvidia RTXA4000 GPU in frames per second (FPS).
@@ 5.3 Results
Figure 4: Visual comparison after two rounds of prune-refine using our and LightGaussian’s methods. Top: bonsai from Mip-NeRF 360. Middle: room from Mip-NeRF 360. Bottom: train from Tanks & Temples. Additional visualizations are presented in Appendix A.8.
@@ 5.3.1 Analysis of PUP 3D-GS Pipeline
The efficacy of our PUP 3D-GS pipeline is demonstrated by Table 1, where we record the mean of each metric over the Mip-NeRF 360 dataset. Pruning $80\%$ of Gaussians from the original 3D-GS model more than doubles its FPS while degrading PSNR, SSIM, and LPIPS. Fine-tuning the pruned model for 5,000 iterations recovers the image quality metrics with only a slight decrease in rendering speed. Pruning $50\%$ of the remaining Gaussians, such that a total of $90\%$ of Gaussians are pruned from the base 3D-GS scene, causes another drop in the image quality metrics. However, we recover most of this degradation by fine-tuning this significantly smaller model for another 5,000 iterations. Rendering speed is more than tripled over the 3D-GS model.
Table 1: Mean PSNR, SSIM, LPIPS, FPS, and point cloud size for the Mip-NeRF 360 dataset at each stage of our PUP 3D-GS pipeline. The operation in each row is applied cumulatively to all of the following rows.
[…]
3D-GS 27.47 0.8123 0.2216 63.88 746.46
+ Prune $80\%$ 21.00 0.7075 0.3075 167.93 149.29
+ Refine 26.97 0.7991 0.2444 148.28 149.29
+ Prune $50\%$ 22.63 0.7021 0.3316 241.97 74.65
+ Refine 26.67 0.7862 0.2719 204.81 74.65
@@ 5.3.2 Comparison with LightGaussian
Figure 3 compares our spatial uncertainty estimate against LightGaussian’s global significance score at different per-round pruning percentages in our two-round prune refine pipeline. After each round of pruning, the models are fine-tuned for $5,000$ iterations for a total of $10,000$ iterations. The per-dataset means of each metric are illustrated as heatmaps. As illustrated by the contour lines, our method consistently outperforms LightGaussian across all metrics and pruning percentage permutations.
We choose to prune $80\%$ then $50\%$ of the Gaussians in the first and second rounds of pruning to optimize image quality and rendering speed at $90\%$ pruning, a significantly higher compression ratio than previous methods. Table 2 compares our $90\%$ pruning results against LightGaussian’s using the per-dataset mean of each metric. We show that our spatial uncertainty estimate outperforms LightGaussian across nearly every metric – the lone exception is PSNR in the Tanks & Temples dataset, where we are outperformed by LightGaussian by $0.36$ points. Our choice of pruning percentages is further ablated in Section 6.2.
Table 2: Comparison of our PUP 3D-GS pipeline against LightGaussian [4]. In both methods, we prune-refine $80\%$ of the Gaussians from the 3D-GS model and then $50\%$ of the remaining Gaussians for a total of $90\%$. The final sizes are identical. PUP 3D-GS increases the rendering speed of 3D-GS by 3.56×3.56$\times$3.56 × on average. Per-scene metrics for each dataset are recorded in Appendix A.4.
@@
MipNeRF-360 3D-GS 27.47 0.8123 0.2216 64.07 746.46
LightGaussian 26.28 0.7622 0.3054 162.12 74.65
Ours 26.67 0.7862 0.2719 204.81 74.65
Tanks & Temples 3D-GS 23.77 0.8458 0.1777 97.86 433.24
LightGaussian 23.08 0.7950 0.2634 329.03 43.33
Ours 22.72 0.8013 0.2441 391.10 43.33
Deep Blending 3D-GS 28.98 0.8816 0.2859 66.79 699.19
LightGaussian 28.51 0.8675 0.3292 234.10 69.92
Ours 28.85 0.8810 0.3015 301.43 69.92
In Figures 1 and 4, we report qualitative results on scenes taken from each dataset. The magnified regions demonstrate that our method retains many fine foreground details that are not preserved by LightGaussian. Error residual visualizations with respect to the ground truth images and original 3D-GS renders are available in Appendix A.10; visualizations of other scenes are provided in Appendix A.8.
Figure 5: Histograms of the distribution of Gaussians over the log of their volumes for the bonsai, room, and train scenes after two rounds of prune-refine. PUP 3D-GS retains smaller Gaussians than LightGaussian, consistent with its higher rendering speed and visual fidelity.
We speculate that the visibility score in LightGaussian’s importance sampling heuristic is biased towards retaining larger Gaussian because they have a higher probability of being visible in more pixels across the training views. In Figure 5, we plot the distributions of the log determinants of the Gaussian covariances over the scenes visualized in Figure 4. The distributions skew more heavily towards larger Gaussians in the scenes pruned with LightGaussian’s pipeline than ours. We address other pruning methods [14, 24, 5] in Appendix A.6.
@@ 5.3.3 Vectree Quantization
We apply Vectree Quantization – a compression method that is orthogonal to pruning – from the LightGaussian framework [4] to further compress our pruned models. Our final models are smaller than the unpruned 3D-GS scenes by an average of 51.70×51.70$\times$51.70 ×, 52.56×52.56$\times$52.56 ×, and 51.06×51.06$\times$51.06 × for the Mip-NeRF 360, Tanks & Temples, and Deep Blending datasets. Detailed results are in Table 3. We also apply our method to EAGLES [6], another orthogonal method, in Appendix A.5.
Table 3: Vectree Quantization comparison with LightGaussian after pruning to $90\%$. PUP 3D-GS outperforms LightGaussian in terms of image quality, rendering speed, and now size.
@@
Mip-NeRF 360 3D-GS 27.47 0.8123 0.2216 64.07 746.46
LightGaussian 24.65 0.7302 0.3341 162.34 14.44
Ours 24.93 0.7584 0.2988 205.97 14.44
Tanks & Temples 3D-GS 23.77 0.8458 0.1777 97.86 433.24
LightGaussian 21.88 0.7679 0.2886 331.47 8.53
Ours 21.61 0.7787 0.2670 389.18 8.49
Deep Blending 3D-GS 28.98 0.8816 0.2859 66.79 699.19
LightGaussian 27.90 0.8586 0.3403 233.28 13.38
Ours 28.24 0.8735 0.3108 300.15 13.30
@@ 6 Ablations
6.1 Multi-Round Pruning
A core component of our pipeline is multiple rounds of prune-refine. In Table 4, we compare the mean image quality metrics and FPS over the Mip-NeRF 360 scenes when we prune using our two step prune-refine method against pruning $90\%$ of Gaussians in a single round with an equivalent number of fine-tuning steps. Two-round pruning produces better measurements across all metrics.
The number of pruning percentage permutations increases exponentially as the number of rounds increases, so we use interpolation to compare two against three rounds of pruning on the Mip-NeRF 360 bicycle scene in Figure 6. Three-round pruning is substantially worse at lower percentages and only slightly better at higher percentages. We choose the simpler two-round approach for our pipeline because both configurations outperform LightGaussian by a much larger margin at $90\%$.
Table 4: One step of pruning $90\%$ of Gaussians from 3D-GS and then fine-tuning for 10,000 iterations, versus pruning $80\%$ then $50\%$ of Gaussians and fine-tuning for 5,000 iterations in each step. In both methods, $90\%$ of Gaussians are pruned cumulatively and the model is fine-tuned for 10,000 total steps. Our multi-round approach outperforms the one-round approach across all metrics.
@@ 6.1 Multi-Round Pruning
Prune $90\%+10$K Fine-tune 26.12 0.7761 0.2807 189.95 74.65
Prune $80\%+50\%$ ($5$K Each) 26.67 0.7862 0.2719 204.81 74.65
6.2 Per-Round Pruning Percentages
We choose to prune $90\%$ of Gaussians, substantially more than previous methods, because preserving visual fidelity at extreme compression ratios is challenging. Figure 7 plots the metrics for each permutation of per-round pruning percentages that results in approximately $90\%$ total pruning. Pruning $80\%$ then $50\%$ of the Gaussians in the first and second rounds optimizes image quality and rendering speed at exactly $90\%$ total pruning. Our method outperforms LightGaussian across all metrics and percentage permutations.
@@ 6.2 Per-Round Pruning Percentages
Another component of our method is restricting our sensitivity pruning score to only the mean and scale parameters of each Gaussian. We choose to exclude rotations, the remaining geometric parameters, because they do not induce a change in 3D geometry when invariances are present. Furthermore, including rotation parameters produces $10\times 10$ Fisher Information matrices that are 2.78×2.78$\times$2.78 × larger than the $6\times 6$ mean and scale parameter-only matrices used in our method, inducing significant additional computational overhead. It is no longer possible to run the pruning pipeline on the Nvidia RTXA4000 GPU used for our other experiments due to memory constraints, necessitating more expensive hardware like the Nvidia RTXA5000.
Table 5 shows that image quality decreases if we include the rotation parameters in our mean and scale parameter sensitivity score. Image quality also decreases if we use the RGB color parameters of the Gaussians to compute our sensitivity score instead. We speculate that the lower performance of the RGB sensitivity score is due to its similarity with LightGaussian’s visibility score – the change in the L2 error over the training images with respect to perturbations in the RGB value of a Gaussian correlates with its visibility across the training views.
Figure 6: Summary results on the MipNeRF-360 bicycle scene for 2×2$\times$2 × and 3×3$\times$3 × prune-refine rounds. We plot the mean and standard deviation of 2×2$\times$2 × for all cumulative pruning percents in Figure 3, computed via cubic interpolation, then do the same for 3×3$\times$3 × with $10\%$ pruning intervals. The dotted red line denotes our target $90\%$ cumulative percentage; per-round results are ablated in Figure 7.
Figure 7: Results for the MipNeRF-360 bicycle scene when pruning approximately $90\%$ of Gaussians with varying per-round percentages in our two-round pipeline. The dotted black line denotes our chosen pruning percentages of $80\%$ then $50\%$; the $90\%$ means from Figure 6 are also shown as dotted lines. Appendix A.3 reports the average metrics across all scenes in the MipNeRF-360 dataset.
Table 5: Results from our two step prune-refine pipeline when computing sensitivity via our score, our score with rotation parameters, and RGB color parameters. Our spatial score outperforms the spatial with rotation and RGB color scores.
Methods Mip-NeRF 360
@@
3D-GS 27.47 0.8123 0.2216 95.59 746.46
Ours 26.67 0.7862 0.2719 204.81 74.65
Ours + Rotation 26.79 0.7861 0.2726 177.30 74.65
RGB Sensitivity 26.40 0.7798 0.2769 195.76 74.65
@@ 7 Limitations
A limitation of our approach is that it assumes that the scene is converged because the Fisher approximation requires a small L1 residual to be accurate. We find that a small residual is preserved even after removing $80\%$ of the Gaussians with prune-refine, allowing for a second round of pruning, as shown in Section 4.4. Another limitation is that the memory requirement for computing the Fisher matrices is proportional to $N\times 36$, where $N$ is the total number of Gaussians. We do not find this size to be prohibitive for our post-hoc pruning pipeline on 16GB Nvidia RTXA4000 or larger GPUs. Appendix A.7 discusses potential directions for future research related to PUP 3D-GS.
@@ 8 Conclusion
In this work, we propose PUP 3D-GS: a new post-hoc pipeline for pruning pretrained 3D-GS models without changing their training pipelines. It uses a novel, mathematically principled approach for choosing which Gaussians to prune by assessing their sensitivity to the reconstruction error over the training views. This sensitivity pruning score is derived from a computationally tractable Fisher approximation of the Hessian over that reconstruction error. We use our sensitivity score to prune Gaussians from the reconstructed scene, then fine-tune the remaining Gaussians in the model over multiple rounds. After pruning $90\%$ of Gaussians, PUP 3D-GS increases average rendering speed by 3.56×3.56$\times$3.56 ×, while retaining more salient foreground information and achieving higher image quality metrics than existing techniques on scenes from the Mip-NeRF 360, Tanks & Temples, and Deep Blending datasets.
@@ Appendix A Appendix
A.1 Fisher Information Derivation: A Bayesian Interpretation of PUP 3D-GS
The Fisher Information is the variance of the score:
$I(\theta)=E_{\theta}[(\nabla_{\theta}\log p(x|\theta))^{2}].$ (11)
Lemma 5.3 from [15] gives that this is equivalent to:
$I(\theta)=E_{\theta}[-\nabla_{\theta}\nabla_{\theta}\log p(x|\theta)],$ (12)
assuming $\log p(x|\theta)$ is twice differentiable and with certain regularity conditions.
To start, we reformulate our $L_{2}$ objective as a log-likelihood:
$-\log p(\mathcal{I}|\Phi,\mathcal{G})=E_{(I,\phi)\sim(\mathcal{I},\Phi)}[(I-I_{\mathcal{G}}(\phi))^{T}(I-I_{\mathcal{G}}(\phi))],$ (13)
where $\mathcal{I}$ is the set of ground truth image, $\Phi$ is the set of their corresponding poses, $\mathcal{G}$ are the model Gaussian parameters, $I_{\mathcal{G}}(\phi)$ is the rendering function for pose $\phi$, and we assume that the error follows a Gaussian distribution.
We can take the Laplace approximation of the log of the posterior distribution over the model parameters $\mathcal{G}$ on the converged scene parameters $\hat{\mathcal{G}}$ as:
$-\log p(\mathcal{G}|\mathcal{I},\Phi)\approx-\log p(\hat{\mathcal{G}}|\mathcal{I},\Phi)+\frac{1}{2}(\mathcal{G}-\hat{\mathcal{G}})H(\hat{\mathcal{G}})(\mathcal{G}-\hat{\mathcal{G}}),$ (14)
where:
$H(\hat{\mathcal{G}})=-\nabla_{\mathcal{G}}\nabla_{\mathcal{G}}\log p(\hat{\mathcal{G}}|\mathcal{I},\Phi).$ (15)
If we assume a uniform prior, then our claimed Fisher Information matrix from Section  4.1 is precisely the Hessian $H(\hat{\mathcal{G}})$ of this posterior.
From this formulation of our Fisher Information matrix, Proposition 3.5 from [12] gives that the log determinant of the Fisher as the entropy of the second order approximation of $p(\mathcal{G}|\mathcal{I},\Phi$) around $\hat{\mathcal{G}}$. If we restrict the posterior to a particular Gaussian’s parameters $\mathcal{G}_{i}$, giving $p(\mathcal{G}_{i}|\mathcal{I},\Phi)$, the log determinant of the block diagonal element corresponding to this Gaussian is a measure of entropy for that particular Gaussian. This interpretation gives our pruning scores as a ranking of the Gaussians by their entropy on this posterior.
A.2 Ablation on Patch Size
We ablate our choice of patch size in Table 6. Notice that, although the $4\times 4$ patches that we use in our experiments produce slightly better image quality metrics, the $2\times 2$ and $8\times 8$ patches also produce similar results.
Table 6: Mean PSNR, SSIM, LPIPS, FPS, and point cloud size for the Mip-NeRF 360 dataset using our sensitivity score computed with $2\times 2$, $4\times 4$ and $8\times 8$ patches.
Methods Mip-NeRF 360
[…]
3D-GS 27.47 0.8123 0.2216 83.88 746.46
$2\times 2$ 26.46 0.7869 0.2742 218.59 74.65
$4\times 4$ (Ours) 26.67 0.7862 0.2719 204.81 74.65
$8\times 8$ 26.53 0.7775 0.2780 189.23 74.65
A.3 Ablation on Per-Round Pruning Percentages
Figure 8 plots the average metrics for each permutation of per-round pruning percentages that results in approximately $90\%$ total pruning across all scenes in the Mip-NeRF 360 dataset. Similar to the bicycle scene evaluated in Figure 7, pruning $80\%$ of Gaussians in the first round and then $50\%$ in the second optimizes image quality and rendering speed at exactly $90\%$ total pruning. Our method outperforms LightGaussian across all metrics and per-round pruning percentage permutations.
Figure 8: The average PSNR, SSIM, LPIPS, and FPS across the MipNeRF-360 dataset scene after pruning approximately $90\%$ of Gaussians with different per-round percentages in our two-round pipeline. The dotted red line denotes our chosen per-round pruning percentages of $80\%$ then $50\%$.
A.4 Scene Evaluations
PSNR, SSIM, LPIPS, and FPS for each scene from the Mip-NeRF 360, Tanks&Temples, and Deep Blending datasets that was used in 3D-GS [11] are recorded in Tables 7, 8, 9, and 10, respectively. Note that the sizes of the pruned scenes in this section are identical because exactly $90\%$ of Gaussians were removed from each of them using our two step prune-refine method. FPS is collected using a Nvidia RTXA4000 GPU.